# Long-lived oscillations of false and true vacuum states in neutral atom systems

### v1

@@ V Deconfinement From Rydberg Tails
The classical model is thus accurate when the kinetic energy term $T_{nn^{\prime}}$ is much less than the potential different $\Delta U_{0}$ in the 1-meson model in Equation 11. In that case, the 1-meson model becomes $U_{0}(n)\phi_{\ell}(n)\approx(\varepsilon_{\ell}/2)\phi_{\ell}(n)$. The largest kinetic energy term has scaling $T_{n,n+2}\sim\Omega^{2}/\Delta_{\mathrm{glob}}$, ignoring smaller effects from $V_{1}$ and $V_{2}$, since expanding a domain wall from size $n$ to $n+2$ involves two Rabi transitions, a second order process, and the intermediate state with 00 sites yields a global detuning penalty, as shown by the term $T_{13}$ in Figure 5. The potential difference from size $n$ to $n+2$ has scaling $\Delta U_{0}\sim\Delta_{\mathrm{loc}}$, again ignoring smaller effects from $V_{1}$ and $V_{2}$. The ratio of the potential to kinetic terms is $\sim\Delta_{\mathrm{loc}}\Delta_{\mathrm{glob}}/\Omega^{2}=\beta\alpha^{2}$, and thus our classical model is an accurate approximation for $\beta\alpha^{2}\gg 1$. The competition is manifested visually in the energy eigenvalues of $\tilde{H}_{0}$ (approximately those of $H$): $T_{n,n+2}$ scales with the energy range, or curvature, of the 1-meson bands across momentum $k$, and $U_{0}(n)$ scales with the energy difference $\omega_{ft}\propto\Delta_{\mathrm{loc}}$ between the false and true vacua. For instance, the case $\beta=\pm 0.1$ gives $\beta\alpha^{2}=1.6$ and is shown in Figure 4i, whereas the case $\beta=\pm 0.01$ gives $\beta\alpha^{2}=0.16$ and is shown in Figure 3c; the case $\beta=\pm 0.1$ is closer to the classical limit, and thus the 1-meson bands are more flat and the vacuum states have larger separation. However, the classical energy model is also approximately valid for larger kinetic energy, since the leading effect of the latter is insensitive to the domain size, i.e. it shifts the 1-meson energies by a term largely independent of the domain size [26].
We can compare our work to that of Lagnese et al. [31], who performed a supplementary study of false and true vacuum oscillations in neutral atom chains with Rydberg excitations. The features we observe for the false vacuum case in Figure 4g for $\beta=0.10$ appear similar to those observed in Ref. [31] Figure S1b for $\beta=0.125$. Furthermore, the features we observe for the true vacuum case in Figure 4h for $\beta=-0.10$ appear similar to those observed in Ref. [31] Figure S1a for $\beta=-0.125$. Lagnese et al. [31] interpreted their observations in analogy with their main results for the ferromagnetic Ising spin chain. However, in our detailed investigations with neutral atom chains that include long-range Rydberg interactions, we found that the peaks have different structure, exhibit different shifting behavior, and arise from different microphysics, all distinct to these antiferromagnetic Rydberg atom systems.
@@ VI Experimental Accessibility
In Section IV, we presented a numerical setup to study false and true vacuum long-lived oscillations, with the following input parameters: $\Omega/2\pi=1.0$ MHz ($R_{b}=9.76$ $\mu$m); $a=5.42$ $\mu$m; $\Delta_{\mathrm{glob}}/2\pi=4.0$ MHz; and $\Delta_{\mathrm{loc}}/2\pi\in[0.0008,0.06]$ MHz.
@@ VII Conclusion and Outlook
In this paper, we studied long-lived oscillations of false and true vacua using neutral atoms with long-range Rydberg interactions. Our work examined setups accessible to near-term neutral atom experiments. We computed the Fourier spectra of quasiparticle oscillations initiated by false and true vacua, and identified spectral features to distinguish between these two cases. In particular, we discovered features distinct to our antiferromagnetic system compared to ferromagnetic Ising chains, and explained them in terms of deconfinement effects from the long-range Rydberg tails.
Our results show that the features of false and true vacuum oscillation spectra contain fingerprints of the underlying confining potential. As a consequence of this dependence on the microscopic details, it is crucial that one examines the underlying microscopic model through simulations before making definite conclusions about the stability of the vacuum. This underscores the importance of nascent quantum simulators in addressing metastability phenomenology.
Our work motivates several further studies. Experiments on 1D neutral atom systems, including analog quantum simulators [35, 36], can probe the oscillation spectra that we observed, motivated by our parameter range (Section IV) and accessibility discussion (Section VI). However, as noted in Section VI, a detailed investigation of the secondary peaks will require intense Rabi pulses to enhance the effective simulation duration or improvements in the coherence time to extend the physical simulation duration. Long-lived oscillations can also be studied on 2D lattices [47, 48]; this could access novel metastability phenomenology that only emerges in 2D [72]. In 2D systems that grow beyond $\mathcal{O}(100)$ sites, classical simulation becomes increasingly intractable as a tool to study false and true vacuum oscillations [73, 74]. Early quantum devices could thus be an exclusive window into the physics of higher-dimensional quantum field theories.

### v2

@@ V Short-Range Meson Repulsion
The classical model is thus accurate when the kinetic energy term $T_{nn^{\prime}}$ is much less than the potential different $\Delta U_{0}$ in the 1-meson model in Equation 11. In that case, the 1-meson model becomes $U_{0}(n)\phi_{\ell}(n)\approx(\varepsilon_{\ell}/2)\phi_{\ell}(n)$. The largest kinetic energy term has scaling $T_{n,n+2}\sim\Omega^{2}/\Delta_{\mathrm{glob}}$, ignoring smaller effects from $V_{1}$ and $V_{2}$, since expanding a domain wall from size $n$ to $n+2$ involves two Rabi transitions, a second order process, and the intermediate state with 00 sites yields a global detuning penalty, as shown by the term $T_{13}$ in Figure 5. The potential difference from size $n$ to $n+2$ has scaling $\Delta U_{0}\sim\Delta_{\mathrm{loc}}$, again ignoring smaller effects from $V_{1}$ and $V_{2}$. The ratio of the potential to kinetic terms is $\sim\Delta_{\mathrm{loc}}\Delta_{\mathrm{glob}}/\Omega^{2}=\beta\alpha^{2}$, and thus our classical model is an accurate approximation for $\beta\alpha^{2}\gg 1$. The competition is manifested visually in the energy eigenvalues of $\tilde{H}_{0}$ (approximately those of $H$): $T_{n,n+2}$ scales with the energy range, or curvature, of the 1-meson bands across momentum $k$, and $U_{0}(n)$ scales with the energy difference $\omega_{mg}\propto\Delta_{\mathrm{loc}}$ between the metastable and ground states. For instance, the case $\beta=\pm 0.1$ gives $\beta\alpha^{2}=1.6$ and is shown in Figure 4i, whereas the case $\beta=\pm 0.01$ gives $\beta\alpha^{2}=0.16$ and is shown in Figure 3c; the case $\beta=\pm 0.1$ is closer to the classical limit, and thus the 1-meson bands are more flat and the ground and metastable states have larger separation. However, the classical energy model is also approximately valid for larger kinetic energy, since the leading effect of the latter is insensitive to the domain size, i.e. it shifts the 1-meson energies by a term largely independent of the domain size [26].
We can compare our work to that of Lagnese et al. [31], who performed a supplementary study of metastable oscillations in neutral atom chains with Rydberg excitations. The features we observe for the metastable case in Figure 4g for $\beta=0.10$ appear similar to those observed in Ref. [31] Figure S1b for $\beta=0.125$. Furthermore, the features we observe for the ground case in Figure 4h for $\beta=-0.10$ appear similar to those observed in Ref. [31] Figure S1a for $\beta=-0.125$. Lagnese et al. [31] interpreted their observations in analogy with their main results for the ferromagnetic Ising spin chain. However, in our detailed investigations with neutral atom chains that include long-range Rydberg interactions, we found that the peaks have different structure, exhibit different shifting behavior, and arise from different microphysics, all distinct to these antiferromagnetic Rydberg atom systems.
@@ VI Experimental Accessibility
In Section IV, we presented a numerical setup to study metastable long-lived oscillations, with the following input parameters: $\Omega/2\pi=1.0$ MHz ($R_{b}=9.76$ $\mu$m); $a=5.42$ $\mu$m; $\Delta_{\mathrm{glob}}/2\pi=4.0$ MHz; and $\Delta_{\mathrm{loc}}/2\pi\in[0.0008,0.06]$ MHz.
@@ VII Conclusion and Outlook
In this paper, we studied long-lived oscillations of metastable and ground states using neutral atoms with long-range Rydberg interactions. Our work examined setups accessible to near-term neutral atom experiments. We computed the Fourier spectra of quasiparticle oscillations initiated by metastable and ground states, and identified spectral features to distinguish between these two cases. In particular, we discovered features distinct to our antiferromagnetic system compared to ferromagnetic Ising chains, and explained them in terms of confinement and short-range domain wall repulsion effects.
Our results show that the features of metastable and ground state oscillation spectra contain fingerprints of the underlying confining potential. As a consequence of this dependence on the microscopic details, it is crucial that one examines the underlying microscopic model through simulations before drawing universal conclusions about metastability. This underscores the importance of nascent quantum simulators in addressing metastability phenomenology.
Our work motivates several further studies. Experiments on 1D neutral atom systems, including analog quantum simulators [36, 37], can probe the oscillation spectra that we observed, motivated by our parameter range (Section IV) and accessibility discussion (Section VI). However, as noted in Section VI, a detailed investigation of the secondary peaks will require intense Rabi pulses to enhance the effective simulation duration or improvements in the coherence time to extend the physical simulation duration. Long-lived oscillations can also be studied on 2D lattices [48, 49]; this could access novel metastability phenomenology that only emerges in 2D [73]. In 2D systems that grow beyond $\mathcal{O}(100)$ sites, classical simulation becomes increasingly intractable as a tool to study metastable oscillations [74, 75]. Early quantum devices could thus be an exclusive window into the physics of higher-dimensional quantum field theories.